An initially motionless test car is accelerated uniformly to 115 km/h in 8.23 seconds before striking a simulated deer. The car is in contact with the faux fawn for 0.575 seconds, after which the car is measured to be traveling at 87.5 km/h. Answers must be to the correct number of sig figs.

a) What is the magnitude of the acceleration of the car before the collision?

b) What is the magnitude of the average acceleration of the car during the collision?

c) What is the magnitude of the average acceleration of the car during the entire test, from when the car first begins moving until the collision is over?
Answer #1

115 Km/h = 31.9444 m/s

so using kinematics, we have

v = u + at

v = at

a = v/t

a = 31.9444 / 8.23

a = 3.8815 m/s2

---------------------------------------------------------------------------------

average acceleration = 24.305 - 31.9444 / 0.575

average acceleration = 13.285 m/s2

--------------------------------------------------------------------------------

average acceleration = 24.305 / 8.23 + 0.575

average acceleration = 2.76 m/s2